OpenStudy (anonymous):

Is csc squaredx + cot squaredx=1 a identity?

OpenStudy (anonymous):

No, \[ \csc^2(x) + \cot^2(x) = \frac{1+\cos^2(x)}{\sin^2(x)}\]

OpenStudy (anonymous):

so is (cscx+cotx) squared=1 an identity? i dont understand how to solve them

OpenStudy (luigi0210):

no, but 1 + cot^2 = csc^2 is an identity

OpenStudy (anonymous):

What do you mean, how to solve them?

OpenStudy (anonymous):

ii dont understand how to see if their identities or not. I have asked my teacher but he doesnt explain anything to me.

OpenStudy (anonymous):

The fundamental identity is \[ \sin^2(x) + \cos^2(x) = 1\] All of the other identities follow from that. Write things in terms of sin and cos and see what happens.

OpenStudy (anonymous):

okay thank you so much! im just trying to get all of my work finished because I will be having a baby tomorrow and wont have time to finish school. I only have 5 quiz's to take but im struggling.

OpenStudy (anonymous):

Well, I wish you the best of luck!!!

OpenStudy (anonymous):

thank you!